OpenStudy (anonymous):

Find the equation of the line through the point (4, 7) that cuts off the least area from the first quadrant. Give your answer using the form below. y - A = B(x - C) A= B= C=

OpenStudy (amistre64):

would that be the area bound by the y-axis, x-axis, and a 45 degree line of negtive slope thru the point (4,7) ?

OpenStudy (anonymous):

I tried everything and all I get was wrong answers

OpenStudy (amistre64):

Q1 is pretty large :)

OpenStudy (amistre64):

what have you tried so far?

OpenStudy (anonymous):

Thank You for trying, I just called my friend, he is willing to show and teach me.

OpenStudy (anonymous):

A=7,B=7/4,C=4 y-7=7/4(x-4) 4(y-7)=7(x-4) 4y-28=7x-28 7x-4y=0